You have decided to open a restaurant in huntington the following table gives the payoffs based on two states of natures favorable and unfavorable demand. You believe that the probability of favorable market is 0.8. You can open any one of three different sizes of restaurant

Open favourable Unfavorable
large restaurant 55000 -30000
Medium restaurant 40000 -5000
Small Restaurant 20000 15000

Find the best choice and provide the necessary analysis to support your answer
Answer #1

here expected value of large Restaurant=0.8*55000+0.2*(-30000)=38000

expected value of medium Restaurant=0.8*40000+0.2*(-5000)=31000

expected value of small Restaurant=0.8*20000+0.2*(15000)=19000

as expected value of large Restaurant is maximum; best choice is Large Restaurant